# DexCap: Scalable and Portable Mocap Data Collection System for Dexterous Manipulation

### v1

@@ Abstract
Imitation learning from human hand motion data presents a promising avenue for imbuing robots with human-like dexterity in real-world manipulation tasks. Despite this potential, substantial challenges persist, particularly with the portability of existing hand motion capture (mocap) systems and the difficulty of translating mocap data into effective control policies. To tackle these issues, we introduce DexCap, a portable hand motion capture system, alongside DexIL, a novel imitation algorithm for training dexterous robot skills directly from human hand mocap data. DexCap offers precise, occlusion-resistant tracking of wrist and finger motions based on SLAM and electromagnetic field together with 3D observations of the environment. Utilizing this rich dataset, DexIL employs inverse kinematics and point cloud-based imitation learning to replicate human actions with robot hands. Beyond learning from human motion, DexCap also offers an optional human-in-the-loop correction mechanism to refine and further improve robot performance. Through extensive evaluation across six dexterous manipulation tasks, our approach not only demonstrates superior performance but also showcases the system’s capability to effectively learn from in-the-wild mocap data, paving the way for future data collection methods for dexterous manipulation. More details can be found at https://dex-cap.github.io
@@ I Introduction
  • DexIL: an imitation learning framework leveraging hand mocap data and point cloud observation for directly learning dexterous manipulation skills.
@@ II-A Dexterous manipulation
Dexterous manipulation has been a long-standing research area in robotics [65, 49, 51, 5, 41], posing significant challenges to planning and control due to the high degrees-of-freedom. The traditional optimal control methods [51, 5, 41] often necessitate simplification of the contacts, which is usually not tenable in more complex tasks. Recently, reinforcement learning has been explored to learn dexterous policies in simulation with minimal assumptions about the task or the environment [28, 11, 10, 23, 85, 59, 37, 31, 12, 57, 82]. The learned policies can solve complex tasks, including in-hand object re-orientatation [28, 11, 85, 59, 37, 57], bimanual manipulation [31, 43], and long-horizon manipulation [12]. However, due to the sim-to-real gap, deploying the learned policy on a real-world robot remains challenging. Imitation learning, on the other hand, focuses on learning directly from real-world demonstration data, which is obtained through either teleportation [27, 3, 2, 24, 86] or human videos [61, 45, 69]. DIME  [3] uses VR to teleoperate a dexterous hand for data collection;  Qin et al. [60] uses a single RGB camera to track hand pose for teleoperation; DexMV [61], DexVIP [45] and VideoDex [69] leverages human video data for learning the motion priors but often require additional training in simulation or real robot teleoperation data. Our work focuses on dexterous imitation learning, which relies on DexCap to collect high-quality hand mocap data grounded in 3D point cloud observation, which can be directly used to train low-level positional control on robots with single or dual dexterous hands.
@@ II-B Hand motion capture system
Human hand mocap is an important technique for applications in computer vision and graphics. Most previous systems are camera-based, IMU-based, or electromagnet(EMF)-based. Camera-based systems utilize monocular camera [89, 50, 55], RGB-D camera [68, 25, 9], VR headset [26], or multi-view camera with markers [76, 16]. However, the quality of hand motion tracking quickly deteriorates in scenarios involving heavy occlusions, which happen frequently in hand-object interactions [16]. Some of these systems also require third-view calibrated cameras which are not portable or scalable. More recently, Inertia Measurement Unit (IMU) has been used for in-the-wild human mocap [32, 36, 84, 77, 80]. Nevertheless, most of them focus on whole-body motion capture and miss fine-grained finger motions. EMF-based mocap gloves are designed for capturing finger motion, which is widely used for dexterous teleoperation [20, 17, 88]. However, the glove does not track the 6-DoF palm poses grounded in the environment and misses visual observations for training robot policies. DexCap is a mocap glove system that is designed to collect data for training visuomotor manipulation policies. Through novel engineering designs, our system stays robust to occlusions, captures fine-grained finger motion, tracks palm poses using SLAM, and records RGB-D images to reconstruct the scene with a wearable camera vest.
@@ II-C Robot learning with human demonstration
Imitation Learning (IL) has enabled robots to successfully perform various manipulation tasks [8, 34, 66, 40, 15, 18, 7, 1]. Traditional IL algorithms such as DMP and PrMP [67, 39, 53, 54] enjoy high learning sample efficiency but are limited in their ability to handle high-dimensional observations. In contrast, recent IL methods built upon deep neural networks can learn policies with raw image observation inputs [48, 19], even for high-degree robot systems with bimanual arms [87, 21]. Despite their effectiveness, one key challenge for imitation learning is how to scale up the training data. Most prior works focus on teleoperation data which is expensive to collect due to the requirement of the robot hardware. More recently, learning from human motion data has started to receive more attention because it allows collecting data without robot hardware [14]. By leveraging human videos [83, 6], hand trajectories [78, 22], promising results have been shown to train policies with less manual human effort. However, most of these human motions are in 2D image space, which fails to directly train 6-DoF manipulation policies in 3D environments and usually requires additional teleoperation data to bridge the gap [83, 6, 78, 22]. Our learning algorithm DexIL, utilizing the high-quality 3D data collected with DexCap, can directly learn dexterous control policies without the need of teleoperation data. Recently, human-in-the-loop correction algorithms have also shown promising results in robot learning [44, 56, 75]. DexCap also provides an optional human-in-the-loop correction mechanism to further improve the performance in more complex tasks.
@@ III Hardware System: DexCap
Figure 3: Details of the robot system. Mirroring the human system, the robot system reuses the same chest cameras and mount. (a) Once the motion is captured by DexCap, it’s retargeted to LEAP hand through discarding pinky finger and IK to match fingertip location. (c) An optional human-in-the-loop correction step can be performed to further refine the motions transferred. Specifically, the human will provide the delta input in real-time when the robot system is carrying out the task. Note the hand T265 is only used at correction time, as the robot arm already knows the exact location of fingers.
@@ III-A DexCap
Tracking finger motions. Our system uses electromagnetic field (EMF) gloves, offering a significant advantage over vision-based finger tracking systems, particularly in the robustness to visual occlusions that frequently occur in hand-object interactions. In our system, finger motions are tracked using Rokoko motion capture gloves as illustrated in Figure 2. Each glove’s fingertip is embedded with a tiny magnetic sensor, while a signal receiver hub is placed on the glove’s dorsal side. The 3D location of each fingertip is measured as the relative 3D translation from the hub to the sensors. In Figure 8 we included a qualitative comparison between our EMF glove system and state-of-the-art vision-based hand-tracking methods across different manipulation scenarios.
@@
Recording 3D observations and calibration. Capturing the data necessary for training robot policies requires not only the tracking of hand movement but also recording observations of the 3D environment as the policy input. As depicted in Figure 2(a), we design a wearable camera vest for this purpose. It incorporates an Intel Realsense L515 RGB-D LiDAR camera, mounted on the top of the chest, to capture the observations during human data collection. The next critical question then becomes how to effectively integrate the tracked hand motion data with the 3D observations. To simplify the calibration process, we designed a 3D-printed camera rack underneath the chest camera mount as illustrated in Figure 2(c). At the beginning of the data collection, all tracking cameras are placed in the rack slots, which secures a constant transformation between the camera frames. Then, we take off the tracking cameras from the rack and insert them into the camera slot attached to each glove. In this way, we can easily transform the hand pose tracking results into the observation frame of the chest camera with the constant initial transformation. To further ensure stable observations amidst human movement, another fisheye tracking camera (marked red in Fig. 2(c)) is mounted under the LiDAR camera, which provides a more robust SLAM performance than the LiDAR camera with its wide field of view. We define the initial pose frame of this tracking camera as the world frame for all stream data. Figure 6 is the visualization of the collected data by transforming the observations into colored point clouds in the world frame alongside the captured hand motions.
@@ IV-B Point cloud-based Diffusion Policy
With the transformed robot’s state $\bm{s}_{t}$, action $\bm{a}_{t}$ and corresponding 3D point cloud observation $\bm{o}_{t}$, we formalize the robot policy learning process as a trajectory generation task. More specifically, an policy model $\pi$, processes the point cloud observations $\bm{o}_{t}$ and the robot’s current proprioception state $\bm{s}_{t}$ into an action trajectory $(\bm{a}_{t},\bm{a}_{t+1},\dots,\bm{a}_{t+d})$ (as in Fig. 4). Given point cloud observation with $N$ points $\bm{o}_{t}$ in $\mathbb{R}^{N\times 3}$, we uniformly down-sample it into $K$ points and concatenate the RGB pixel color corresponding to each point into the final policy input in $\mathbb{R}^{K\times 6}$. To bridge the visual gap between human hands and the robot’s hand, we use forward kinematics to transform the links of the robot model with the proprioception state $\bm{s}_{t}$ and merge the point clouds of the transformed links into the observation $\bm{o}_{t}$. During training, we also use data augmentation over the inputs by applying random 2D translations to the point clouds and motion trajectories within the robot’s operational space.
One challenge of learning dexterous robot policies, especially for bimanual dexterous robots, is handling the large dimensional action outputs. In our setup, the action output includes two 7-DoF robot arms and two 16-DoF dexterous hands for $d$ steps, which forms a high-dimensional regression problem. Similar challenges have also been studied in image generation tasks, which aim to regress all pixel values in a high-resolution frame. Recently, diffusion model [73, 30], with its step-by-step diffusion process, has shown success in modeling complex data distributions with high-dimensional data. For robotics, Diffusion Policy [13] follows the same idea and formalizes the control problem into an action generation task. Thus we use Diffusion Policy as the action decoder, where we empirically find it outperforms traditional MLP-based architecture for learning dexterous robot policies.
@@ IV-C Human-in-the-loop correction
Figure 7: Objects used in the Packaging task
Figure 8: Compare with SOTA vision-based method. We demonstrate that our motion capture data significantly surpasses the quality of the state-of-the-art vision-based pose estimation method HaMeR [55] when facing similar occlusion challenges.
@@ V-A Experiment setups
We utilize three data types: (1) DexCap data capturing human hand motion within the robot’s operational space, (2) in-the-wild DexCap data from outside lab environments, and (3) human-in-the-loop correction data for adjusting robot actions or enabling teleoperation to correct errors, collected using a foot pedal. Data were initially recorded at 60Hz and then downsampled to 20Hz to match the robot’s control speed, except for correction data, which was collected directly at 20Hz. For data collection, we gathered 30 minutes of DexCap data across the first three tasks, resulting in 201, 129, and 82 demos respectively. An hour of in-the-wild DexCap data provided 96 demos for Packaging. Scissor Cutting and Tea Preparing tasks each received an hour of DexCap data, yielding 104 and 55 demos respectively.
@@
Each model variant is tested for 20 trials in each task with randomized initial placements. The task success rate is reported in Table. I II III. For the multi-object Packaging task, each object is tested with 5 trials - 6 trained objects (30 total trials) and 9 unseen objects (45 total trials), as is shown in Figure. 7.
@@ V-B Results
DexCap delivers high-quality 3D mocap data ($\mathbf{\mathcal{Q}1}$). Figure 6 showcases DexCap’s ability to capture detailed hand motion in 3D, aligning human actions with object point clouds across all views, such as in Plate wiping and Scissor cutting tasks (blue columns). The retargeted robot hand motions, depicted in the yellow columns, demonstrate precise alignment in the same 3D space. In Figure 8, we compare DexCap with the state-of-the-art vision-based hand pose estimation method HaMeR [55], observing their performance from similar viewpoints. We find that the vision-based approach is vulnerable to self-occlusion, particularly when the fingers are obscured. As depicted in Figure 8, HaMeR struggles in instances of significant occlusion, either failing to detect the hand (as seen in the second column) or inaccurately estimating fingertip positions (noted in the first, third, and fourth columns). In contrast, DexCap demonstrates good robustness under these conditions. Beyond the challenge of occlusion, most vision-based methods rely on 2D hand estimation, predicated on learning from 2D image projection losses. Consequently, these methods are inherently limited in their ability to discern the precise 3D hand positioning, as they are trained based on presumed, fixed camera intrinsic parameters, which do not necessarily match the actual camera used for experiments. These results highlight DexCap’s capability in collecting high-quality dexterous manipulation data. In addition, DexCap data is fast to collect. As is illustrated in Figure. 9, DexCap is about three times faster than the traditional teleoperation method in data collection throughput and is close to the same level of natural human motion.
@@
Human-in-the-loop correction greatly help when DexCap data is insufficient ($\mathbf{\mathcal{Q}5}$). Figure. 10 illustrates two types of human-in-the-loop correction mode with DexCap. Users can switch between the two modes by stepping on the foot pedal and the whole trajectory is stored and used for fine-tuning the policy. The last three columns of Table II showcase the effectiveness of using human-in-the-loop correction together with policy fine-tuning to improve the model performance. With just 30 human correction trials during policy rollout, the fine-tuned policy with image inputs (DP-img-mask) achieves a 33% improvement in the full task success rate for trained objects. This significant boost is mainly because the human correction data is collected using a fixed camera - the same setup used for the evaluations. This result further supports our conclusion: image-based approaches are more effective in learning with fixed third-view cameras compared to the in-the-wild scenarios with moving cameras. Human corrections also result in a 10% improvement in our approach that utilizes point cloud inputs. However, we’ve observed that fine-tuning with human corrections has a minor effect on the results for unseen objects, primarily due to the limited amount of correction data (30 trials in total).
@@
VI Conclusion
We present DexCap, a portable hand motion capture system, and DexIL, an imitation algorithm enabling robots to learn dexterous manipulation directly from human mocap data. DexCap, designed to overcome occlusions, capture fine-grained 3D hand motion, record RGB-D observations, and allow data collection outside the lab. DexIL applies this data to teach robots complex dexterous manipulation tasks, with an optional human-in-the-loop correction mechanism to further improve performance. Demonstrating proficiency in tasks like scissor cutting and tea preparation, DexCap and DexIL significantly advance robotic dexterity. In future works, we aim to leverage the internet-scale human hand motion data to pretrain the policy. Hierarchical policy architectures like [78] could also show effects in such scenarios. We hope DexCap can pave the path for future research on scaling up dexterous manipulation data with portable devices. All hardware designs and code are open-sourced.

### v2

@@ Abstract
Imitation learning from human hand motion data presents a promising avenue for imbuing robots with human-like dexterity in real-world manipulation tasks. Despite this potential, substantial challenges persist, particularly with the portability of existing hand motion capture (mocap) systems and the complexity of translating mocap data into effective robotic policies. To tackle these issues, we introduce DexCap, a portable hand motion capture system, alongside DexIL, a novel imitation algorithm for training dexterous robot skills directly from human hand mocap data. DexCap offers precise, occlusion-resistant tracking of wrist and finger motions based on SLAM and electromagnetic field together with 3D observations of the environment. Utilizing this rich dataset, DexIL employs inverse kinematics and point cloud-based imitation learning to seamlessly replicate human actions with robot hands. Beyond direct learning from human motion, DexCap also offers an optional human-in-the-loop correction mechanism during policy rollouts to refine and further improve task performance. Through extensive evaluation across six challenging dexterous manipulation tasks, our approach not only demonstrates superior performance but also showcases the system’s capability to effectively learn from in-the-wild mocap data, paving the way for future data collection methods in the pursuit of human-level robot dexterity.
@@ I Introduction
  • DexIL: an imitation learning framework leveraging hand mocap data for directly learning dexterous manipulation skills from human hand motions.
@@ II-A Dexterous manipulation
Dexterous manipulation has been a long-standing research area in robotics [15, 16, 17, 18, 19], posing significant challenges to planning and control due to the high degrees-of-freedom. The traditional optimal control methods [17, 18, 19] often necessitate simplification of the contacts, which is usually not tenable in more complex tasks. Recently, reinforcement learning has been explored to learn dexterous policies in simulation with minimal assumptions about the task or the environment [20, 2, 21, 22, 23, 24, 25, 26, 27, 28, 29]. The learned policies can solve complex tasks, including in-hand object re-orientatation [20, 2, 23, 24, 25, 28], bimanual manipulation [26, 30], and long-horizon manipulation [22, 27]. However, due to the sim-to-real gap, deploying the learned policy on a real-world robot remains challenging. Imitation learning, on the other hand, focuses on learning directly from real-world demonstration data, which is obtained through either teleportation [1, 31, 32, 6] or human videos [33, 34, 3]. DIME [31] uses VR to teleoperate a dexterous hand for data collection;  Qin et al. [35] uses an RGB camera to track hand pose for teleoperation; DexTransfer [36] uses human mocap data to guide dexterous grasping; DexMV [33], DexVIP [34] and VideoDex [3] leverages human video data for learning the motion priors but often require additional training in simulation or real robot teleoperation data. Our work focuses on dexterous imitation learning, which relies on DexCap to collect high-quality hand mocap data grounded in 3D point cloud observation, which can be directly used to train low-level positional control on robots with single or dual hands.
@@ II-B Hand motion capture system
Human hand mocap is an important technique for applications in computer vision and graphics. Most previous systems are camera-based, IMU-based, or electromagnet(EMF)-based. Camera-based systems utilize monocular camera [37, 38, 39], RGB-D camera [40, 41, 42], VR headset [43], or multi-view camera with markers [44, 45]. However, the quality of hand motion tracking quickly deteriorates in scenarios involving heavy occlusions, which happen frequently in hand-object interactions. Some of these systems also require third-view calibrated cameras which are not portable or scalable. More recently, Inertia Measurement Unit (IMU) has been used for in-the-wild human mocap [46, 47, 48, 49, 50]. Nevertheless, most of them focus on whole-body motion capture and miss fine-grained finger motions. EMF-based mocap gloves are designed for capturing finger motion, which is widely used for dexterous teleoperation [51, 52, 53]. However, the glove does not track the 6-DoF palm poses grounded in the environment and misses visual observations for training robot policies. DexCap is a mocap glove system that is designed to collect data for training visuomotor manipulation policies. Through novel engineering designs, our system stays robust to occlusions, captures fine-grained finger motion, tracks palm poses using SLAM, and records RGB-D images to reconstruct the scene with a wearable camera vest.
@@ II-C Robot learning with human demonstration
Imitation Learning (IL) has enabled robots to successfully perform various manipulation tasks [54, 55, 4, 56, 57, 58, 59, 60]. Traditional IL algorithms such as DMP and PrMP [61, 62, 63, 64] enjoy high learning sample efficiency but are limited in their ability to handle high-dimensional observations. In contrast, recent IL methods built upon deep neural networks can learn policies with raw image observation inputs [65, 66], even for high-degree robot systems with bimanual arms [67, 68]. Despite their effectiveness, one key challenge for imitation learning is how to scale up the training data. Prior works focus on teleoperation data  [69, 66, 70, 71, 72, 73, 74, 75, 76, 77] which is expensive to collect due to the requirement of the robot hardware. More recently, learning from human motion data has started to receive more attention because it allows collecting data without robot hardware [78]. By leveraging human videos [79, 11], hand trajectories [80, 10, 81, 82], promising results have been shown to train policies with less manual human effort. However, these human motions are in 2D image space  [83, 84, 80], which fails to directly train 6-DoF manipulation policies in 3D environments and usually requires additional teleoperation data to bridge the gap  [79, 11, 10]. Recently, human-in-the-loop correction algorithms have also shown promising results in robot learning [85, 86, 87]. Our DexCap provides tracking of 6-DoF hand poses together with finger motions grounded in 3D point cloud observations, which is portable for data collection without a robot. Based on the data collected with DexCap, we introduce DexIL which is a point cloud-based imitation learning algorithm for learning fine-grained dexterous manipulation policies, with an optional human-in-the-loop correction step for more challenging tasks.
II-D Portable data collection systems for manipulation
Recently advancements in low-cost hand-held grippers have shown promising results in collecting robot manipulation data without robot hardware  [88, 89, 90, 91, 92, 93, 94]. All of these systems are designed and used for the parallel-gripper data collection process, while in this work we aim to collect multi-finger hand motion data for dexterous manipulation tasks (e.g., using scissors and unscrewing bottle caps).
@@ III Hardware System: DexCap
Figure 3: Details of the robot system. Mirroring the human system, the robot system reuses the same chest cameras and mount. (a) Once the motion is captured by DexCap, it’s retargeted to LEAP hand through discarding pinky finger and IK to match fingertip location. (c) An optional human-in-the-loop correction step can be performed to further refine the motions transferred. Specifically, the human will provide the delta input in real time when the robot system is carrying out the task. Note the hand T265 is only used at correction time, as the robot arm already knows the exact location of fingers.
@@ III-A DexCap
Tracking finger motions. Our system uses electromagnetic field (EMF) gloves, offering a significant advantage over vision-based finger tracking systems, particularly in the robustness to visual occlusions that frequently occur in hand-object interactions. In our system, finger motions are tracked using Rokoko motion capture gloves as illustrated in Figure 2. Each glove’s fingertip is embedded with a tiny magnetic sensor, while a signal receiver hub is placed on the glove’s dorsal side. The 3D location of each fingertip is measured as the relative 3D translation from the hub to the sensors. In appendix we included a qualitative comparison between our EMF glove system and state-of-the-art vision-based hand-tracking methods across different manipulation scenarios.
@@
Recording 3D observations and calibration. Capturing the data necessary for training robot policies requires not only the tracking of hand movement but also recording observations of the 3D environment as the policy input. As depicted in Figure 2(a), we design a wearable camera vest for this purpose. It incorporates an Intel Realsense L515 RGB-D LiDAR camera, mounted on the top of the chest, to capture the observations during human data collection. The next critical question then becomes how to effectively integrate the tracked hand motion data with the 3D observations. To simplify the calibration process, we designed a 3D-printed camera rack underneath the chest camera mount as illustrated in Figure 2(c). At the beginning of the data collection, all tracking cameras are placed in the rack slots, which secures a constant transformation between the camera frames. Then, we take off the tracking cameras from the rack and insert them into the camera slot attached to each glove. In this way, we can easily transform the hand pose tracking results into the observation frame of the chest camera with the constant initial transformation. The full calibration process is demonstrated in Appendix Figure  13 and supplementary videos, which takes around 10 seconds. To further ensure stable observations amidst human movement, another fisheye tracking camera (marked red in Fig. 2(c)) is mounted under the LiDAR camera, which provides a more robust SLAM performance than the LiDAR camera with its wide field of view. We define the initial pose frame of this tracking camera as the world frame for all stream data. Figure 6 is the visualization of the collected data by transforming the observations into colored point clouds in the world frame alongside the captured hand motions.
@@ IV-B Point cloud-based diffusion policy
With the transformed robot’s state $\bm{s}_{t}$, action $\bm{a}_{t}$ and corresponding 3D point cloud observation $\bm{o}_{t}$, we formalize the robot policy learning process as a trajectory generation task. More specifically, a policy model $\pi$, processes the point cloud observations $\bm{o}_{t}$ and the robot’s current proprioception state $\bm{s}_{t}$ into an action trajectory $(\bm{a}_{t},\bm{a}_{t+1},\dots,\bm{a}_{t+d})$ (as in Fig. 4). Given point cloud observation with $N$ points $\bm{o}_{t}$ in $\mathbb{R}^{N\times 3}$, we uniformly down-sample it into $K$ points and concatenate the RGB pixel color corresponding to each point into the final policy input in $\mathbb{R}^{K\times 6}$. To bridge the visual gap between human hands and the robot’s hand, we use forward kinematics to transform the links of the robot model with the proprioception state $\bm{s}_{t}$ and merge the point clouds of the transformed links into the observation $\bm{o}_{t}$. During training, we also use data augmentation over the inputs by applying random 2D translations to the point clouds and motion trajectories within the robot’s operational space.
One challenge of learning dexterous robot policies, especially for bimanual dexterous robots, is handling the large dimensional action outputs. In our setup, the action output includes two 7-DoF robot arms and two 16-DoF dexterous hands for $d$ steps, which forms a high-dimensional regression problem. Similar challenges have also been studied in image generation tasks, which aim to regress all pixel values in a high-resolution frame. Recently, diffusion model [99, 100], with its step-by-step diffusion process, has shown success in modeling complex data distributions with high-dimensional data. For robotics, diffusion policy [14] follows the same idea and formalizes the control problem into an action generation task. Thus we use a diffusion policy as the action decoder, where we empirically find it outperforms traditional MLP-based architecture for learning dexterous robot policies.
@@ IV-C Human-in-the-loop correction
Figure 7: Compare with vision-based method. We demonstrate that motion capture gloves provide more stable hand pose estimation results compared to vision-based methods and are not affected by visual occlusion.
@@ V-A Experiment setups
We utilize two data types: (1) DexCap data capturing human hand motion (In-the-wild data refers to a mixture of data collected in more than 10 scenes) and (2) human-in-the-loop correction data for adjusting robot actions or enabling teleoperation to correct errors, collected using a foot pedal. Data were initially recorded at 60Hz and then downsampled to 20Hz to match the robot’s control speed, except for correction data, which was collected directly at 20Hz. For data collection, we gathered 30 minutes of DexCap data across the first three tasks, resulting in 251, 179, and 102 demos respectively. An hour of in-the-wild DexCap data provided 104 demos for Packaging. Scissor Cutting and Tea Preparing tasks each received an hour of DexCap data, yielding 96 and 55 demos respectively.
@@
Each model variant is tested for 20 trials in each task with randomized initial placements. The task success rate is reported in Table I II III. For the multi-object Packaging task, each object is tested with 5 trials - 6 trained objects (30 total trials) and 9 unseen objects (45 total trials).
@@ V-B Results
DexCap delivers high-quality 3D mocap data ($\mathbf{\mathcal{Q}1}$). Figure 6 showcases DexCap’s ability to capture detailed hand motion in 3D, aligning human actions with object point clouds across all views, such as in Plate wiping and Scissor cutting tasks (blue columns). The retargeted robot hand motions, depicted in the yellow columns, demonstrate precise alignment in the same 3D space. In Figure 7, we compare DexCap with the state-of-the-art vision-based hand pose estimation method HaMeR [39], observing their performance from similar viewpoints. We find that the vision-based approach is vulnerable to self-occlusion, particularly when the fingers are obscured. As depicted in Figure 7, HaMeR struggles in instances of significant occlusion, either failing to detect the hand (as seen in the second column) or inaccurately estimating fingertip positions (noted in the first, third, and fourth columns). In contrast, DexCap demonstrates good robustness under these conditions. Beyond the challenge of occlusion, most vision-based methods rely on 2D hand estimation, predicated on learning from 2D image projection losses. Consequently, these methods are inherently limited in their ability to discern the precise 3D hand positioning, as they are trained based on presumed, fixed camera intrinsic parameters, which do not necessarily match the actual camera used for experiments. In Figure 8, we showcase the data collection throughput of DexCap, which is three times faster than traditional teleoperation.
@@
Figure 10: Objects used in the Packaging task
Human-in-the-loop correction greatly help when DexCap data is insufficient ($\mathbf{\mathcal{Q}5}$). Figure 9 illustrates two types of human-in-the-loop correction mode with DexCap. Users can switch between the two modes by stepping on the foot pedal and the whole trajectory is stored and used for fine-tuning the policy. The last three columns of Table II showcase the effectiveness of using human-in-the-loop correction together with policy fine-tuning to improve the model performance. With just 30 human correction trials during policy rollout, the fine-tuned policy with image inputs (DP-img-mask) achieves a 33% improvement in the full task success rate for trained objects. This significant boost is mainly because the human correction data is collected using a fixed camera - the same setup used for the evaluations. This result further supports our conclusion: image-based approaches are more effective in learning with fixed third-view cameras compared to the in-the-wild scenarios with moving cameras. Human corrections also result in a 10% improvement in our approach that utilizes point cloud inputs. However, we’ve observed that fine-tuning with human corrections has a minor effect on the results for unseen objects, primarily due to the limited amount of correction data (30 trials in total).
@@
VI Conclusion and Limitations
We present DexCap, a portable hand motion capture system, and DexIL, an imitation algorithm enabling robots to learn dexterous manipulation directly from human mocap data. DexCap, designed to overcome occlusions, capture fine-grained 3D hand motion, record RGB-D observations, and allow data collection outside the lab. DexIL applies this data to teach robots complex dexterous manipulation tasks, with an optional human-in-the-loop correction mechanism to further improve performance. Demonstrating proficiency in tasks like scissor cutting and tea preparation, DexCap and DexIL significantly advance robotic dexterity. We hope DexCap can pave the path for future research on scaling up dexterous manipulation data with portable devices. All hardware designs and code will be open-source.
While DexCap collects high-quality mocap data in-the-wild for learning challenging dexterous manipulation tasks, it has several limitations that need future research: (1) The system’s power consumption currently restricts the collection time to be at most 40 minutes. Future improvements will focus on enhancing power efficiency to extend the collection time. (2) Our learning algorithm DexIL utilizes fingertip inverse kinematics to retarget human hand motion to various robotic hands. However, the size difference between human and robotic hands (with some robotic fingers being thicker) can make some tasks difficult to perform, such as playing the piano. Future developments will aim to integrate advancements in robotic hand design to minimize these size differences and fully demonstrate the system’s potential. (3) Current DexCap collects only 3D observations and motion capture data, lacking force sensing. One promising direction we plan to explore involves the use of conformal tactile textiles, as introduced in [107], to gather tactile information during data collection.